# Predicting Grain Boundary Segregation in Magnesium Alloys: An Atomistically Informed Machine Learning Approach

### v1

@@ Abstract
Grain boundary (GB) segregation in magnesium (Mg) substantially influences its mechanical properties and performance. Atomic-scale modelling, typically using ab-initio or semi-empirical approaches, has mainly focused on GB segregation at highly symmetric GBs in Mg alloys, often failing to capture the diversity of local atomic environments and segregation energies, resulting in inaccurate structure-property predictions. This study employs atomistic simulations and machine learning models to systematically investigate the segregation behavior of common solute elements in polycrystalline Mg at both ground state and finite temperatures. The machine learning models accurately predict segregation thermodynamics by incorporating energetic and structural descriptors. We found that segregation energy and vibrational free energy follow skew-normal distributions, with hydrostatic stress, an indicator of excess free volume, emerging as an important factor influencing segregation tendency. The local atomic environment’s flexibility, quantified by flexibility volume, is also crucial in predicting GB segregation. Comparing the grain boundary solute concentrations calculated via the Langmuir–McLean isotherm with experimental data, we identified a pronounced segregation tendency for Nd, highlighting its potential for GB engineering in Mg alloys. This work demonstrates the powerful synergy of atomistic simulations and machine learning, paving the way for designing advanced lightweight Mg alloys with tailored properties.
@@ I Introduction
Due to their low density and superior mechanical properties, magnesium (Mg) alloys are promising lightweight structural materials for enhancing sustainable, low-carbon automotive applications [1, 2]. However, their widespread commercial adoption is hindered by substantial processing barriers, notably their limited formability, and ductility due to anisotropic plasticity and strong crystallographic texture [3, 4]. In response, recent years have seen significant efforts to improve the strength and ductility of Mg alloys. These advancements involve introducing alloying elements to reduce the disparity in critical resolved shear stresses between basal and non-basal dislocations [5, 6] and weakening the strong basal texture via grain boundary (GB) segregation [7, 8].
Solute segregation at GBs has been widely reported to influence GB energy and mobility, subsequently affecting the texture development of Mg alloys [9, 10]. Understanding the atomistic origins of GB segregation phenomena is crucial for designing Mg alloys with specifically tailored properties. Nie et al. [11] reported periodic substitutional segregation of Gd and Zn solutes at twin boundaries in Mg alloys using high-resolution scanning transmission electron microscopy and attributed this periodic solute decoration to strain energy minimization using the density functional theory (DFT) calculations. The pinning effect of this ordered solute segregation on twin boundaries leads to annealing strengthens of these Mg alloys. Similar atomic-resolution electron microscopic approaches have been carried out on other highly symmetric tilt GBs in Mg alloys, where characteristic solute decoration patterns were similarly observed [12, 13, 14].
Additional experimental efforts have been carried out to quantify the solute concentrations at GBs in Mg alloys [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Pei et al. [30] measured varying Nd solute concentrations between 2 and 5 at.%, across six different GBs in Mg alloys using atom probe tomography (APT). The observed inhomogeneous segregation behavior within the GB plane, which was rationalized through correlated atomistic simulations on general GBs with the same misorientations and GB planes, suggests that it stems from local atomic arrangements within the GBs rather than macroscopic characteristics. Using APT, Mouhib et al. [31] characterized the concentrations of Ca and Gd solutes at GBs in Mg alloys, both with and without adding Zn. The selective formation of specific texture components in the ternary Mg-Zn-M alloys (M = Ca or Gd) was correlated to the synergistic effects of Zn addition on solute clustering and binding via DFT.
@@
Atomic-scale simulations via ab-initio or semi-empirical approaches have mostly focused on GB segregation at highly symmetric GBs in Mg alloys, which may not sufficiently capture the broad spectra of local atomic environments and segregation energies at GBs, leading to inaccurate predictions in structure-property relationships [40]. Although Wagih’s work [35] covers a wide range of solute elements in polycrystalline Mg, the differing interatomic potentials with different Mg-Mg interactions make comparisons difficult, particularly since the embedded-atom method (EAM) potentials [41] were not developed for modeling hexagonal close-packed Mg. Additionally, most solute elements studied are rarely used in Mg-based solid solutions for practical, lightweight applications. Therefore, no experimental data are available for comparison. To date, there has been no comprehensive investigation of the GB segregation spectrum across various common solute elements at both absolute zero and finite temperatures in polycrystalline Mg. In this study, we systematically investigated GB segregation of six solute elements—Nd, Ca, Y, Li, Al, and Zn—commonly found in Mg alloys within a Mg polycrystal using atomistic simulations. Apart from determining the segregation energy spectra at 0 K, the vibrational entropy contribution at finite temperatures was taken into account using the harmonic approximation (HA) approach. The predicted GB solute concentrations were directly compared with experimental data from the literature, covering relevant bulk concentrations and temperature regimes. Furthermore, ML approaches were employed to predict both the segregation energy and the segregation free energy, with the goal of analyzing the essential structural and energetic parameters that provide predictive power for the pure Mg polycrystal.
@@ II.1 Atomistic simulations
Atomistic simulations in this study were performed using the MD software package LAMMPS [42]. Interatomic interactions were modeled by the modified embedded atom method (MEAM) potentials developed by Lee et al. for Mg–M (M = Nd [43], Ca [44], Y [44], Li [45], Al [46] and Zn [47]) systems, where the Mg–Mg interaction [46] is identical. These MEAM potentials have shown good performance in predicting substitutional segregation energies at highly symmetric GBs in Mg [33].
The polycrystalline Mg structure was constructed via Voronoi tessellation using Atomsk [48]. The periodic simulation box, measuring 15 nm in dimensions, contains 12 grains and approximately 145,000 atoms, with 26 % of these sites characterized as GB sites according to the adaptive common neighbor analysis (CNA) [49]. Following initial energy minimization, the sample was heated to 500 K over 50 ps and annealed at 500 K for 20 ps before being quenched back to 0 K over 50 ps using the isothermal-isobaric ensemble with the Nosé-Hoover thermostat and barostat [50]. Subsequently, the FIRE algorithm [51, 52] was applied to relax the structure with a force tolerance of $10^{-8}$ eV/Å.
The flexibility volume ($V_{\text{flex}}$) [53] of a GB site in the polycrystal was calculated following Equation 1:
@@
where $\Omega_{i}$ is the atomic volume of the $i$th atom according to the Voronoi tessellation. The vibrational mean-squared displacement ($MSD_{\text{vib},i}$) of the $i$th atom was calculated using $\left[x_{i}(t_{0})-x_{i}(t_{0}+\Delta t)\right]^{2}$, where $x_{i}(t_{0})$ is the equilibrium position of the $i$th atom and $x_{i}(t_{0}+\Delta t)$ is its position after 5000 MD steps (timestep is 1 fs) under the microcanonical ensemble at 300 K. The overall MSD of the polycrystalline system is flat with time and contains the vibrational but not the diffusional contribution, see Figure S1. The averaged vibrational mean-squared displacement $MSD_{\text{vib},i}\right>$ was calculated by averaging over 100 independent MD runs with different initial velocity distributions.
@@ II.2 Segregation energy and vibrational entropy contribution
The segregation energy $\Delta E_{\text{seg}}$ at 0 K for a substitutional solute atom M at a GB site was calculated using Equation 2:
$\Delta E_{\text{seg}}=(E_{\text{bulk}}-E_{\text{bulk,M}})-(E_{\text{GB}}-E_{\text{GB,M}}),$ (2)
where $E_{\text{bulk}}$ represents the energy of a bulk system, $E_{\text{bulk,M}}$ denotes the energy of the bulk system with a host atom replaced by an M solute, $E_{\text{GB}}$ is the energy of a system containing a GB, and $E_{\text{GB,M}}$ stands for the energy of the system with a solute atom M occupying the GB site. The structures were relaxed using the FIRE algorithm with a force tolerance of $10^{-8}$ eV/Å. A negative segregation energy indicates favorable solute segregation.
The segregation free energy $\Delta F_{\text{seg}}$ of a substitutional solute in a solid at finite temperatures $T$ was computed by considering the contribution of entropy change (excess vibrational entropy) $\Delta S_{\text{seg}}$, as given by Equation 3:
@@
with the segregation vibrational free energy $\Delta F_{\text{seg}}^{\text{vib}}$ calculated using Equation 5:
$\Delta F_{\text{seg}}^{\text{vib}}=(F_{\text{bulk}}^{\text{vib}}-F_{\text{bulk,M}}^{\text{vib}})-(F_{\text{GB}}^{\text{vib}}-F_{\text{GB,M}}^{\text{vib}}),$ (5)
where $F_{\text{bulk}}^{\text{vib}}$ represents the vibrational free energy of a bulk system, $F_{\text{bulk,M}}^{\text{vib}}$ denotes the vibrational free energy of the bulk system with a host atom replaced by an M solute, $F_{\text{GB}}^{\text{vib}}$ is the vibrational free energy of a system containing a GB, and $F_{\text{GB,M}}^{\text{vib}}$ stands for the vibrational free energy of the system with a solute atom M occupying the GB site.
Figure 1: (a) The vibrational heat capacities at constant volume $C_{\text{v}}$ and pressure $C_{\text{p}}$ of bulk Mg (containing 5 $\times$ 5 $\times$ 5 orthogonal unit cells) as functions of temperature, calculated by the full harmonic approximation (HA) method using the MEAM potential. For comparison, results from density functional perturbation theory (DFPT) [54] and experimental data [55] are presented. (b) The distribution of $F^{\text{vib}}$ of all GB sites in the Mg polycrystal at 600 K calculated based on HA ($r_{\text{HA}}$=14 Å, bin size: 1×$\times 10^{-4}$ eV/atom). A skew-normal function (dashed line) was fitted to the $F^{\text{vib}}$ distribution. The dashed line indicates the $F^{\text{vib}}$ value for bulk Mg.
@@
where $M_{i}$ is the atomic mass of atom $i$, $u_{i\alpha}$ is the displacement of the $i$ atom in the direction $\alpha$, $U$ is the potential energy of the system, and $\Phi_{i\alpha j\beta}^{0}$ is the force constant of atom $i$ in the direction $\alpha$ when atom $j$ is displaced in the $\beta$ direction. Equation 6 can be rewritten by looking for solutions of the form $u_{i\alpha}(t)=1/\sqrt{M_{i}}\tilde{u}_{i\alpha}\exp(2\pi i\nu t)$:
@@ II.3 Machine learning prediction
In this study, we applied ML to predict per-site segregation energy $\Delta E_{\text{seg}}$, segregation vibrational free energy $\Delta F_{\text{seg}}^{\text{vib}}$, and segregation free energy $\Delta F_{\text{seg}}$ at 600 K in Mg alloys using datasets representing six alloying elements: Nd, Ca, Y, Li, Al and Zn. Each dataset contained 38,114 GB sites and 16 energetic and structural descriptors (ESD), including potential energy, six stress tensors ($\sigma_{ij}$), hydrostatic stress ($\sigma_{h}$), coordination number, centrosymmetric parameter (CSP), structural features obtained from Voronoi tessellation (atomic volume, cavity radius, and surface area of the polyhedra), vibrational free energy $F^{\text{vib}}$ at 600 K, vibrational MSD ($MSD_{\text{vib}}$), and flexibility volume ($V_{\text{flex}}$). We preprocessed the data by selecting relevant features, excluding the target variables, and performing correlation analysis using heatmaps. In addition to ESD, we generated the fingerprint of the local atomic environment at GB sites using SOAP [58, 59]. The SOAP approach models the Gaussian particle density functions for all neighboring atoms in the local atomic environment using a combination of radial basis functions and spherical harmonics. The parameters controlling the SOAP descriptor vector include the cutoff radius ($r_{\text{max}}$ = 6 Å), the maximum number of radial basis functions ($n_{\text{max}}$ = 12), the degree of spherical harmonics ($l_{\text{max}}$ = 12), and the Gaussian function width ($\sigma_{\text{at}}$ = 1 Å). In total, the SOAP vector comprises 1015 features. The SOAP descriptors were combined with the ESD features to enhance prediction accuracy. For the prediction of $\Delta F_{\text{seg}}^{\text{vib}}$, $\Delta E_{\text{seg}}$ was included as an input variable along with other features.
@@
The CatBoostRegressor model [60], known for its ability to handle categorical data efficiently and its robust performance with minimal hyperparameter tuning, was chosen for this task. Each dataset was split into training, testing, and validation sets (70-20-10 ratio). The split allocated 70% of the data to training, 20% to testing, and 10% to validation. The hyperparameter space was explored using GridSearchCV with a predefined parameter grid, which included various combinations of iterations, learning rate, tree depth, L2 regularization, and bagging temperature, to identify the optimal settings that maximize the $R^{2}$ score. The model was trained with the optimized hyperparameters using 1000 iterations, a learning rate of 0.03, a tree depth of 8, an L2 regularization term of 1, and a bagging temperature of 0.5 to ensure convergence, see Figure S2(d). 5-fold cross-validation was performed on the training set to evaluate the model’s robustness and stability. Model performance was evaluated on both the testing and validation sets using mean absolute error (MAE) and $R^{2}$ scores. We visualized the results through scatter plots of true versus predicted values with Gaussian Kernel Density Estimation (KDE) to highlight prediction densities, histograms comparing the distributions of true and predicted per-site energies, and feature importance plots.
@@
To benchmark the performance of the ML models, we evaluated GradientBoostingRegressor [61], XGBoostRegressor [62], and ElasticNetCV [63] models implemented in scikit-learn [64] in addition to CatBoostRegressor on the same datasets consisting of the ESD features using the default hyperparameters. The evaluation metrics demonstrated that all models showed similar behavior, but CatBoostRegressor slightly outperformed the others, achieving the lowest MAE values and the highest $R^{2}$ scores across all datasets in predicting $\Delta E_{\text{seg}}$ in Mg alloys, making it the most reliable model for this application. The benchmarking results are compiled in Table S1 and Figure S3, providing a comprehensive comparison of model performances.

### v2

@@ Abstract
Grain boundary (GB) segregation substantially influences the mechanical properties and performance of magnesium (Mg). Atomic-scale modeling, typically using ab-initio or semi-empirical approaches, has mainly focused on GB segregation at highly symmetric GBs in Mg alloys, often failing to capture the diversity of local atomic environments and segregation energies, resulting in inaccurate structure-property predictions. This study employs atomistic simulations and machine learning models to systematically investigate the segregation behavior of common solute elements in polycrystalline Mg at both 0 K and finite temperatures. The machine learning models accurately predict segregation thermodynamics by incorporating energetic and structural descriptors. We found that segregation energy and vibrational free energy follow skew-normal distributions, with hydrostatic stress, an indicator of excess free volume, emerging as an important factor influencing segregation tendency. The local atomic environment’s flexibility, quantified by flexibility volume, is also crucial in predicting GB segregation. Comparing the grain boundary solute concentrations calculated via the Langmuir–McLean isotherm with experimental data, we identified a pronounced segregation tendency for Nd, highlighting its potential for GB engineering in Mg alloys. This work demonstrates the powerful synergy of atomistic simulations and machine learning, paving the way for designing advanced lightweight Mg alloys with tailored properties.
@@ I Introduction
Magnesium (Mg) alloys are lightweight structural materials with promising applications in sustainable, low-carbon industries, despite challenges with limited formability and ductility due to strong crystallographic texture [1, 2, 3, 4, 5, 6, 7]. Addressing these issues involves alloying to balance critical resolved shear stresses [8, 9, 10] and leveraging solute grain boundary (GB) segregation to influence texture [11, 12, 13]. Solute segregation at GBs has been widely reported to influence GB energy and mobility, subsequently affecting the texture development of Mg alloys [14, 15]. Understanding the atomistic origins of GB segregation phenomena is crucial for designing Mg alloys with specifically tailored properties.
Nie et al. [16] reported periodic substitutional segregation of Gd and Zn solutes at twin boundaries in Mg alloys using high-resolution scanning transmission electron microscopy and attributed this periodic solute decoration to strain energy minimization using the density functional theory (DFT) calculations. The pinning effect of this ordered solute segregation on twin boundaries leads to the annealing strengthening of these Mg alloys. Similar atomic-resolution electron microscopic approaches have been carried out on other highly symmetric tilt GBs in Mg alloys, where characteristic solute decoration patterns were similarly observed [17, 18, 19]. Extensive experimental efforts have been made to quantify solute concentrations at GBs in Mg alloys. This was driven by the fact that certain alloying elements (rare earth, Y, and Ca) exhibit a pronounced tendency to segregate at these boundaries, even at low concentrations within the Mg matrix, involving intricate atomic-scale interactions [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Pei et al. [35] measured varying Nd solute concentrations between 2 and 5 at.%, across six different GBs in Mg alloys using atom probe tomography (APT). The observed inhomogeneous segregation behavior within the GB plane, which was rationalized through correlated atomistic simulations on general GBs with the same misorientations and GB planes, suggests that it stems from local atomic arrangements within the GBs rather than macroscopic characteristics. Using APT, Mouhib et al. [36] characterized the concentrations of Ca and Gd solutes at GBs in Mg alloys, both with and without adding Zn. The selective formation of specific texture components in the ternary Mg-Zn-X alloys (X= Ca or Gd) was correlated to the synergistic effects of Zn addition on solute clustering and binding via DFT.
@@
Atomic-scale simulations via ab-initio or semi-empirical approaches have mostly focused on GB segregation at highly symmetric GBs in Mg alloys, which may not sufficiently capture the broad spectra of local atomic environments and segregation thermodynamic quantities at GBs, leading to inaccurate predictions in structure-property relationships [45]. Although Wagih’s work [40] covers a wide range of solute elements in polycrystalline Mg, the differing interatomic potentials with different Mg-Mg interactions make comparisons difficult, particularly since the embedded-atom method (EAM) potentials [46] were not developed for modeling hexagonal close-packed Mg. Additionally, most solute elements studied are rarely used in newly developed Mg-based solid solutions for practical, lightweight applications. Therefore, no experimental data are available for comparison with the theoretical and computational approaches. To date, there has been no comprehensive investigation of the GB segregation spectrum across various common solute elements in polycrystalline Mg at both absolute zero and finite temperatures. In this study, we systematically investigated GB segregation of six solute elements—Nd, Ca, Y, Li, Al, and Zn—commonly used in wrought Mg alloy sheets and extrusions. Our investigations employed atomistic simulations within a magnesium polycrystal framework comprising a broad variety of site types of the full GB space. In addition to determining the segregation energy spectra at 0 K, we also accounted for the vibrational entropy contribution at finite temperatures using the harmonic approximation (HA) approach. The predicted GB solute concentrations were directly compared with experimental data from the literature, covering relevant bulk concentrations and temperature regimes. Furthermore, ML approaches were employed in combination with atomistic modeling to predict both the segregation energy and the segregation free energy, with the goal of providing rapid predictions using the essential structural and energetic parameters within the pure Mg polycrystal at a reduced computational cost.
@@ II.1 Atomistic simulations
Atomistic simulations in this study were performed using the molecular dynamics (MD) software package LAMMPS [47]. Interatomic interactions were modeled by the modified embedded atom method (MEAM) potentials developed by Lee et al. for Mg–X (X= Nd [48], Ca [49], Y [49], Li [50], Al [51] and Zn [52]) systems, where the Mg–Mg interaction [51] is identical. These MEAM potentials have shown good performance in predicting substitutional segregation energies at highly symmetric GBs in Mg [38].
The polycrystalline Mg structure with a random texture was constructed via Voronoi tessellation using Atomsk [53]. The periodic simulation box, measuring 15 nm in dimensions, contains 12 grains and approximately 145,000 atoms, with 26 % of these sites characterized as GB sites according to the adaptive common neighbor analysis (CNA) [54]. Following initial energy minimization, the sample was heated to 500 K over 50 ps and annealed at 500 K for 20 ps before being quenched back to 0 K over 50 ps using the isothermal-isobaric ensemble with the Nosé-Hoover thermostat and barostat [55]. Subsequently, the FIRE algorithm [56, 57] was applied to relax the structure with a force tolerance of $10^{-8}$ eV/Å.
The flexibility volume ($V_{\text{flex}}$[58] of a GB site in the polycrystal was calculated following Equation (1):
@@
where $\Omega_{i}$ is the atomic volume of the $i$th atom according to the Voronoi tessellation. The vibrational mean-squared displacement ($MSD_{\text{vib},i}$) of the $i$th atom was calculated using $\left[x_{i}(t_{0})-x_{i}(t_{0}+\Delta t)\right]^{2}$, where $x_{i}(t_{0})$ is the equilibrium position of the $i$th atom and $x_{i}(t_{0}+\Delta t)$ is its position after 5000 MD steps (timestep is 1 fs). The system was initialized by assigning velocities corresponding to an initial temperature of 600 K and then simulated under the microcanonical (NVE) ensemble. Due to the equipartition of the potential energy and kinetic energy, the energy fluctuates around half of the initially assigned temperature value and effectively maintains the temperature near 300 K during the simulation. The overall MSD of the polycrystalline system is flat with time and contains the vibrational but not the diffusional contribution, see Figure S1 (Supplementary Material). The averaged vibrational mean-squared displacement $MSD_{\text{vib},i}\right>$ was calculated by averaging over 100 independent MD runs with different initial velocity distributions.
@@ II.2 Segregation energy and vibrational entropy contribution
The segregation energy $\Delta E_{\text{seg}}$ at 0 K for a substitutional solute atom X at a GB site was calculated using Equation (2):
$\Delta E_{\text{seg}}=(E_{\text{bulk}}-E_{\text{bulk,X}})-(E_{\text{GB}}-E_{\text{GB,X}}),$ (2)
where $E_{\text{bulk}}$ represents the energy of a bulk system, $E_{\text{bulk,X}}$ denotes the energy of the bulk system with a host atom replaced by an X solute, $E_{\text{GB}}$ is the energy of a system containing a GB, and $E_{\text{GB,X}}$ stands for the energy of the system with a solute atom X occupying the GB site. The structures were relaxed using the FIRE algorithm with a force tolerance of $10^{-8}$ eV/Å. A negative segregation energy indicates favorable solute segregation.
The segregation free energy $\Delta F_{\text{seg}}$ of a substitutional solute in a solid at finite temperatures $T$ was computed by considering the contribution of entropy change (excess vibrational entropy) $\Delta S_{\text{seg}}$, as given by Equation (3):
@@
with the segregation vibrational free energy $\Delta F_{\text{seg}}^{\text{vib}}$ calculated using Equation (5):
$\Delta F_{\text{seg}}^{\text{vib}}=(F_{\text{bulk}}^{\text{vib}}-F_{\text{bulk,X}}^{\text{vib}})-(F_{\text{GB}}^{\text{vib}}-F_{\text{GB,X}}^{\text{vib}}),$ (5)
where $F_{\text{bulk}}^{\text{vib}}$ represents the vibrational free energy of a bulk system, $F_{\text{bulk,X}}^{\text{vib}}$ denotes the vibrational free energy of the bulk system with a host atom replaced by an X solute, $F_{\text{GB}}^{\text{vib}}$ is the vibrational free energy of a system containing a GB, and $F_{\text{GB,X}}^{\text{vib}}$ stands for the vibrational free energy of the system with a solute atom X occupying the GB site.
@@
where $M_{i}$ is the atomic mass of atom $i$, $u_{i\alpha}$ is the displacement of the $i$ atom in the direction $\alpha$, $U$ is the potential energy of the system, and $\Phi_{i\alpha j\beta}^{0}$ is the force constant of atom $i$ in the direction $\alpha$ when atom $j$ is displaced in the $\beta$ direction. Equation (6) can be rewritten by looking for solutions of the form $u_{i\alpha}(t)=1/\sqrt{M_{i}}\tilde{u}_{i\alpha}\exp(2\pi i\nu t)$:
@@ II.3 Machine learning prediction
In this study, we applied ML to predict per-site segregation energy $\Delta E_{\text{seg}}$, segregation vibrational free energy $\Delta F_{\text{seg}}^{\text{vib}}$, and segregation free energy $\Delta F_{\text{seg}}$ at 600 K in binary Mg alloys using datasets representing six alloying elements: Nd, Ca, Y, Li, Al and Zn. Each dataset contained 38,114 GB sites and 16 energetic and structural descriptors (ESD), including potential energy, six stress tensors ($\sigma_{ij}$), hydrostatic stress ($\sigma_{h}$), coordination number, centrosymmetric parameter (CSP), structural features obtained from Voronoi tessellation (atomic volume, cavity radius, and surface area of the polyhedra), vibrational free energy $F^{\text{vib}}$ at 600 K, vibrational MSD ($MSD_{\text{vib}}$), and flexibility volume ($V_{\text{flex}}$). We preprocessed the data by selecting relevant features, excluding the target variables, and performing correlation analysis using heatmaps. In addition to ESD, we generated the fingerprint of the local atomic environment at GB sites using SOAP [61, 62]. The SOAP approach models the Gaussian particle density functions for all neighboring atoms in the local atomic environment using a combination of radial basis functions and spherical harmonics. The parameters controlling the SOAP descriptor vector include the cutoff radius ($r_{\text{max}}$ = 6 Å), the maximum number of radial basis functions ($n_{\text{max}}$ = 12), the degree of spherical harmonics ($l_{\text{max}}$ = 12), and the Gaussian function width ($\sigma_{\text{at}}$ = 1 Å). In total, the SOAP vector for each GB site comprises 1015 features. The SOAP descriptors were combined with the ESD features to enhance prediction accuracy. For the prediction of $\Delta F_{\text{seg}}^{\text{vib}}$, $\Delta E_{\text{seg}}$ was included as an input variable along with other features.
@@
The CatBoostRegressor model [63], known for its ability to handle categorical data efficiently and its robust performance with minimal hyperparameter tuning, was chosen for this task. Each dataset was split into training, testing, and validation sets (70-20-10 ratio). The split in the learning algorithm allocated 70% of the data to training, 20% to testing, and 10% to validation. The hyperparameter space was explored using GridSearchCV with a predefined parameter grid, which included various combinations of iterations, learning rate, tree depth, L2 regularization, and bagging temperature, to identify the optimal settings that maximize the $R^{2}$ score. The model was trained with the optimized hyperparameters using 1000 iterations, a learning rate of 0.03, a tree depth of 8, an L2 regularization term of 1, and a bagging temperature of 0.5 to ensure convergence, see Figure S3(d). 5-fold cross-validation was performed on the training set to evaluate the model’s robustness and stability. Model performance was evaluated on both the testing and validation sets using mean absolute error (MAE) and $R^{2}$ scores. We visualized the results through scatter plots of true versus predicted values with Gaussian Kernel Density Estimation (KDE) to highlight prediction densities, histograms comparing the distributions of true and predicted per-site energies, and feature importance plots.
@@
To assess the performance of the proposed ML model, we evaluated GradientBoostingRegressor [64], XGBoostRegressor [65], and ElasticNetCV [66] models implemented in scikit-learn [67] in addition to CatBoostRegressor on the same datasets consisting of the ESD features using the default hyperparameters. The evaluation metrics demonstrated that all models showed similar behavior, but CatBoostRegressor slightly outperformed the others, achieving the lowest MAE values and the highest $R^{2}$ scores across all datasets in predicting $\Delta E_{\text{seg}}$ in the investigated Mg alloys, making it the most reliable model for this application. The benchmarking results are compiled in Table S1 and Figure S3, providing a comprehensive comparison of model performances.